OpenStudy (lena772):

A husband and wife attend the local recreation center. The husband pays a flat rate of $125 for one year, shown as f(x). The wife pays $6.00 per visit, shown as g(x). Which function shows the correct combination of these two functions to represent the total cost to them of attending the recreation center in one year, shown as h(x)?

OpenStudy (lena772):

f(x) = 125x, g(x) = 6, h(x) = 125x + 6 f(x) = 125, g(x) = 6, h(x) = 6 + 125 f(x) = 125x, g(x) = 6x, h(x) = 125x + 6x f(x) = 125, g(x) = 6x, h(x) = 6x + 125

OpenStudy (lena772):

I think it's the second or last

OpenStudy (shamil98):

125 for one year so that's not changing. but 6 dollars per visit so g(x) = 6x and f(x) = 125

OpenStudy (lena772):

because 125 it's a set rate and those are the two options that has it that way

OpenStudy (lena772):

k so the last one :)

OpenStudy (lena772):

thank you!